# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v3

@@ 4 Simulations
    Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
Finally, we compare our bias correction method to the Jackknife approach by Fernández-Val and Weidner (2016), who consider a probit model in their simulation. We consider a similar setup, using the same settings as in (4.1), except that the disturbance term $\varepsilon_{it}$ follows a standard normal distribution. Table 4 displays the estimation bias and RMSE of the uncorrected MLEs, our bias-corrected MLEs, and the Jackknife bias-corrected estimators, both for the static ($\theta_{0}=\beta_{0}$) and dynamic ($\theta_{0}=(\rho_{0},\beta_{0})^{\prime}$) probit models. In the static model, $\widehat{\beta}_{L}$ is our bias-corrected MLE with $\tau=0$, and $\widehat{\beta}_{J}$ represents the Jackknife-corrected MLE. In the dynamic model, $(\widehat{\beta}_{L},\widehat{\rho}_{L})^{\prime}$ is our bias-corrected MLE with $\tau=1$, and $(\widehat{\beta}_{J},\widehat{\rho}_{J})^{\prime}$ is the Jackknife-corrected MLE. Table 4 shows that our bias-corrected MLEs outperform both the uncorrected and Jackknife bias-corrected estimators, with consistently smaller RMSEs across all DGPs. The bias correction by our method is significantly greater than that of the Jackknife approach, especially when the sample size is small. For example, in the static model with $(N,T)=(30,30)$, the bias of our corrected MLE is only $0.048$, while the bias of the Jackknife estimator is $-1.228,$ over twenty-five times larger.
@@
$\widehat{\beta}$ $0.144$ $0.197$ $0.061$ $0.081$ $0.041$ $0.052$
$\widehat{\beta}_{L}$ $0.048$ $0.123$ $0.011$ $0.051$ $0.007$ $0.031$
$\widehat{\beta}_{J}$ $-1.228$ $5.930$ $-0.022$ $0.058$ $-0.007$ $0.032$
@@
$\widehat{\beta}$ $0.172$ $0.262$ $0.069$ $0.088$ $0.042$ $0.054$
$\widehat{\beta}_{L}$ $0.072$ $0.207$ $0.017$ $0.053$ $0.007$ $0.032$
$\widehat{\beta}_{J}$ $-1.339$ $3.873$ $-0.029$ $0.101$ $-0.009$ $0.034$
$\widehat{\rho}$ $-0.050$ $0.122$ $-0.028$ $0.056$ $-0.017$ $0.037$
$\widehat{\rho}_{L}$ $-0.025$ $0.104$ $-0.014$ $0.049$ $-0.007$ $0.032$
$\widehat{\rho}_{J}$ $0.003$ $0.115$ $-0.001$ $0.050$ $0.000$ $0.033$
@@
    Notes: $(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\beta_{0}=0$ for the static model. Accordingly, $(\widehat{\beta},\widehat{\rho})^{\prime}$ (or $\widehat{\beta}$) denotes the uncorrected MLE; $(\widehat{\beta}_{L},\widehat{\rho}_{L})^{\prime}$ (or $\widehat{\beta}_{L}$) is our bias-corrected MLE with truncation parameter $\tau=1$ for the dynamic model or $\tau=0$ for the static model; and $(\widehat{\beta}_{J},\widehat{\rho}_{J})^{\prime}$ (or $\widehat{\beta}_{J}$) represents the Jackknife bias-corrected estimators. The number of replications is $1000$.
@@ Acknowledgement
We would like to express our gratitude to valuable comments and suggestions from Geert Dhaene, Jinyong Hahn, Chen Hsiao, Whitney Newey, and Jun Yu. Our research assistant Yuchi Liu helped us with various tasks for the empirical analysis. Leng’s research is partially supported by the National Natural Science Foundation of China Grant 72001183. Sun acknowledges the support from the National Natural Science Foundation of China under Grant Number 72203032.
@@ A Proofs
A.1 Lemmas
Proof of Lemma 1. Note that ${H}(\theta)=D\big{[}\partial_{\phi\phi^{\prime}}l(\theta,\phi)\mid_{\phi=\phi(\theta)}\big{]}D^{\prime},$ where $D$ is a $(N+T-2)K\times(N+T)K$ matix. Since $K$ is fixed, we assume without loss of generality that $K=1.$ We express $H(\theta)$ as:
$H(\theta)=\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right),&H_{\breve{\alpha}\breve{\gamma}^{\prime}}\left(\theta\right)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}\left(\theta\right),&H_{\breve{\gamma}\breve{\gamma}^{\prime}}\left(\theta\right)\end{pmatrix}.$
Consider any element $h_{i,j}^{\breve{\alpha}}(\theta)$ of $N\cdot H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta),$ where $i,j=1,\ldots,N-1.$ Under Assumptions 1 and 1, by applying Davydov’s inequality, it follows that
$\sup_{i,j}\sup_{\theta\in\Theta}\text{Var}\big{(}h_{i,j}^{\breve{\alpha}}(\theta)\big{)}=o(1)$
as $N$ and $T\rightarrow\infty.$ Consequently, by the Markov inequality,
$\sup_{i,j}\sup_{\theta\in\Theta}\big{\{}h_{i,j}^{\breve{\alpha}}(\theta)-\mathbb{E}[h_{i,j}^{\breve{\alpha}}(\theta)]\big{\}}=o_{\mathbb{P}}(1).$
Let $h_{t,s}^{\breve{\gamma}}(\theta)$ denote the $(t,s)$-th element of $T\cdot H_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)$ for $t,s=1,\ldots,T-1.$ By conditional cross-sectional independence, we also have
$\sup_{t,s}\sup_{\theta\in\Theta}\big{\{}h_{t,s}^{\breve{\gamma}}(\theta)-\mathbb{E}[h_{t,s}^{\breve{\gamma}}(\theta)]\big{\}}=o_{\mathbb{P}}(1).$
Moreover, for each $h_{i,t}^{\breve{\alpha}\breve{\gamma}}(\theta),$ the $(i,t)$-th element of $NT\cdot H_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta),$ with $i=1,\ldots,N-1$ and $t=1,\ldots,T-1,$ by Assumption 1 we have that
$\sup_{i,t}\sup_{\theta\in\Theta}\big{\{}h_{i,t}^{\breve{\alpha}\breve{\gamma}}(\theta)-\mathbb{E}[h_{i,t}^{\breve{\alpha}\breve{\gamma}}(\theta)]\big{\}}=o_{\mathbb{P}}(1).$
Therefore,
$N\left\|H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)-\mathbb{E}[H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),\quad T\left\|H_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)-\mathbb{E}[H_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),$
and
$NT\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)-\mathbb{E}[H_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),$
uniformly for $\theta\in\Theta,$ where $\|A\|_{\max}$ denotes the max norm of a matrix $A$. Since $\mathbb{E}[s(\theta)]=0,$ $s(\theta)=s(\theta)-\mathbb{E}[s(\theta)]=:(\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta),\widetilde{s}_{\breve{\gamma}}^{\prime}(\theta))^{\prime}.$ We now decompose the quadratic form on the right hand side of (2.7):
$s^{\prime}(\theta)[-{H}(\theta)]^{-1}s\left(\theta\right)$
@@ A.1 Lemmas
$=$ $-\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+\widetilde{s}_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)+2\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}\left(\theta\right)\big{]}(1+o_{\mathbb{P}}(1)).$ (A.1)
By matrix inversion in block form, it follows that
$\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}=O(1),$
@@
$\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\right\|_{\max}=O(T),\quad\sup_{\theta\in\Theta}\left\|H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\right\|_{\max}=O(N).$
In the following, we show
$\sup_{\theta\in\Theta}\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{]}=O\big{(}1/(NT)\big{)},$ (A.2)
and
@@
$\sup_{\theta\in\Theta}\,\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\alpha}}(\theta)\big{]}=O(1/T),\quad\sup_{\theta\in\Theta}\,\mathbb{E}\big{[}\widetilde{s}_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\gamma}}(\theta)\big{]}=O(1/N).$ (A.3)

### v4

@@ 4 Simulation
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
Next, we conduct the same experiment under Design 2, where we generate $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$, keeping everything else intact as in (4.1). We increase the volatility of $\alpha_{k,i}^{0}$, $\gamma_{k,t}^{0}$, $X_{it}$, and $X_{i0}$ to induce larger biases and size distortions. This time, we present the results for the probit model and leave that for the logit model in Appendix B.1.
Table 4 presents the comaprison of the bias and RMSE of the corrected and the uncorrected estimators for the probit model, under Design 2. First of all, the probit estimator here is much more biased than the previous logit counterparts. For the static model with $(N,T)=(30,30)$, $\widehat{\beta}$ contains a bias of $0.943$. Recall that the true value is only $\beta_{0}=0.5$, implying that the bias is almost $200\%$. Our corrected estimator $\widehat{\beta}_{L}$ successfully reduces this bias to $0.328$, which is one third of the uncorrected and roughly $60\%$ biased. As the sample sizes increase, the bias systematically becomes smaller both for the uncorrected and corrected estimators, with dramatic improvements observed from the corrected ones both for the stiatic and dynamic models. We indeed observe that the RMSE of $\widehat{\beta}_{L}$ is twice as large as that of $\widehat{\beta}$ for the static model under $N,T=(30,30)$. However, it is the only incidence here and we attribute this to some sort of numerical coincidence [^8] Below, when we compare our method with the Jackknife under the same setting, it will be seen that the RMSE of our corrected estimator is indeed smaller than the uncorrected. See also Footnote 9 below.. It can be checked that the RMSE of our corrected estimator is no worse the uncorrected one in every other case.
Table 4: Comparisons of Uncorrected and Corrected Probit Estimates, Design 2
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.943$ $2.272$ $0.151$ $0.252$ $0.073$ $0.081$
$\widehat{\beta}_{L}$ $0.328$ $5.170$ $0.041$ $0.062$ $0.017$ $0.040$
Dynamic Model
$\widehat{\beta}$ $0.909$ $1.842$ $0.166$ $0.248$ $0.084$ $0.100$
$\widehat{\beta}_{L}^{(1)}$ $0.207$ $0.396$ $0.048$ $0.078$ $0.022$ $0.042$
$\widehat{\beta}_{L}^{(2)}$ $0.202$ $0.282$ $0.049$ $0.080$ $0.022$ $0.042$
$\widehat{\rho}$ $-0.030$ $0.162$ $-0.016$ $0.073$ $-0.011$ $0.045$
$\widehat{\rho}_{L}^{(1)}$ $-0.015$ $0.140$ $-0.005$ $0.066$ $-0.004$ $0.044$
$\widehat{\rho}_{L}^{(2)}$ $-0.009$ $0.141$ $-0.001$ $0.067$ $-0.001$ $0.045$
[…]
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$. For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$. The number of replications is $1000$.
Table 5 below reports the empirical size of the LR test for the probit model under Design 2. As before, the null hypothesis is $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model and $H_{0}:\beta_{0}=0.5$ for the static. We set the truncation parameter $\tau=0$ for the static model and $\tau=1$ for the dynamic model. Here the size distortion is much more fatal than in the logit case above. For the static model with $(N,T)=(30,30)$, the uncorrected LR test produces an empirical size of $77\%$. Even under $(N,T)=(90,90)$, the empirical size is still $72\%$, indicating a prohibitively high risk of committing a Type-I error. On the other hand, the $L(\theta)$-based LR test delivers a $14\%$ empirical size only, which is almost a $60\%$ decrease in magnitude. Similar performance is observed in other cases. The results here demonstrate that our method can reduce the type-I error risks by a major margin. Indeed, our test still over-rejects under $(N,T)=(90,90)$. This is anticipated, because the IPP under Design 2 is much more drastic and $(N,T)=(90,90)$ may not be sufficient for the test statistic to be well-approximated by $\chi^{2}$. However, it is clear that, as $(N,T)$ increase, the empirical size does become smaller. We have tried additional small-scale simulations (not reported though) with even larger $(N,T)$ and found that the empirical size still decreases with $(N,T)$. We similarly report the results from the LM and the Wald test in Appendix B.2. We report the empirical power under Design 2 in the Appendix B, since they do not add too much value in the discussion here. A quick summary for those results is that our procedure produces tests with stronger power in Design 2.
Table 5: Empirical Size of LR Tests Based on Corrected and Uncorrected Likelihoods for Logit Model, Design 2
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $77\%$ $39\%$ $71\%$ $37\%$
$(60,60)$ $76\%$ $20\%$ $66\%$ $19\%$
$(90,90)$ $72\%$ $14\%$ $63\%$ $14\%$
[…]
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level of $5\%$. The number of replications is $1000$.
Finally, we compare our method with the Jackknife of Fernández-Val and Weidner (2016). We consider the same probit model and continue the same setting and design as above. Table 6 displays the estimation bias and the RMSE of the uncorrected estimators $\{\widehat{\rho},\widehat{\beta}\}$, our corrected estimator $\{\widehat{\rho}_{L},\widehat{\beta}_{L}\}$, and the Jackknife-corrected estimator $\{\widehat{\rho}_{J},\widehat{\beta}_{J}\}$ for the dynamic model. For the static model, only $\widehat{\beta}$, $\widehat{\beta}_{L}$, and $\widehat{\beta}_{J}$ apply and are reported. We set the truncation parameter $\tau=0$ for static model and $\tau=1$ for dynamic model. Across all pairs of $(N,T)$, the Jackknife-corrected $\widehat{\beta}_{J}$ consistently exhibits larger bias and RMSE than our $\widehat{\beta}_{L}$. Even worse, $\widehat{\beta}_{J}$ shows a larger bias and RMSE than the uncorrected $\widehat{\beta}$ across all pairs of $(N,T)$. For instance, with $(N,T)=(30,30)$ and the static model, $\widehat{\beta}_{J}$ shows a bias of $-4.929$, which is roughly $900\%$ biased (in magnitude) relative to the true value $\beta_{0}=0.5$. Observe that the uncorrected estimator $\widehat{\beta}$ only shows a roughly $170\%$ bias relative to the true value. Our corrected $\widehat{\beta}_{L}$ performs the best among the three, showing a $36\%$ bias [^9] Also notice that the RMSE of $\widehat{\beta}_{L}$ in Table 6 is much smaller than that of $\widehat{\beta}_{L}$ in Table 4, where both tables are essentially under the same setting. This motivates our attribution of the large RMSE in Table 4 to numerical coincidence.. To summarize, we find that the Jackknife is not a suitable candidate for DN-HP models. Although it does reduce the bias for the homogeneous coefficient for large $N,T$, the bias the Jackknife causes to the heterogeneous parameter is dangerously large, especially in small sample.
@@
$\widehat{\beta}$ $0.846$ $2.168$ $0.151$ $0.261$ $0.074$ $0.082$
$\widehat{\beta}_{L}$ $0.182$ $0.339$ $0.041$ $0.068$ $0.019$ $0.033$
$\widehat{\beta}_{J}$ $-4.929$ $14.383$ $-0.504$ $0.981$ $-0.105$ $0.214$
@@
$\widehat{\beta}$ $1.040$ $2.343$ $0.184$ $0.287$ $0.080$ $0.090$
$\widehat{\beta}_{L}$ $0.209$ $0.332$ $0.052$ $0.082$ $0.020$ $0.049$
$\widehat{\beta}_{J}$ $-3.352$ $7.708$ $-0.558$ $1.125$ $-0.118$ $0.217$
$\widehat{\rho}$ $-0.046$ $0.169$ $-0.022$ $0.072$ $-0.014$ $0.046$
$\widehat{\rho}_{L}$ $-0.028$ $0.145$ $-0.011$ $0.067$ $-0.009$ $0.044$
$\widehat{\rho}_{J}$ $-0.029$ $0.169$ $-0.006$ $0.068$ $-0.002$ $0.044$
@@
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The true value is $(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\beta_{0}=0$ for the static model. Accordingly, $(\widehat{\beta},\widehat{\rho})^{\prime}$ (or $\widehat{\beta}$) denotes the uncorrected estimator; $(\widehat{\beta}_{L},\widehat{\rho}_{L})^{\prime}$ (or $\widehat{\beta}_{L}$) is our bias-corrected MLE with truncation parameter $\tau=1$ for the dynamic model or $\tau=0$ for the static model; and $(\widehat{\beta}_{J},\widehat{\rho}_{J})^{\prime}$ (or $\widehat{\beta}_{J}$) represents the Jackknife bias-corrected estimators. The number of replications is $1000$.
@@ 5 Empirical Analysis
6 Final Remarks
In this paper, we propose a likelihood-based analytical bias correction procedure for a general class of dynamic nonlinear heterogeneous parameter models subject to the incidental parameter problem. We give the analytical form of a corrected likelihood and show that such a corrected likelihood delivers point estimates that are asymptotically unbiased and test statistics that are asymptotically $\chi^{2}$-distributed. Simulation studies and empirical analyses support our claims.
Several issues are, however, not formally addressed in this paper. We conjecture, but have not formally investigated, that the average partial effects associated these models are asymptotically unbiased. Specifically, we conjecture that the rate of convergence of the average partial effects are only $O(1/\sqrt{T})$ or $O(1/\sqrt{N})$, slower than the potential bias, indicating that bias corrections may have no benefit in the limit. This result is formally shown by Fernández-Val and Weidner (2016) for models with additive fixed effects in the intercept. Second, we have not applied our procedure to general M-estimation models. We intuitively believe that our approach applies to them as long as the objective function is differentiable, only with minor modifications regarding the information matrix equality. We would like to dedicate these extensions to future research.
@@ Acknowledgement
We would like to express our gratitude to valuable comments and suggestions from Geert Dhaene, Yanqin Fan, Jinyong Hahn, Chen Hsiao, Whitney Newey, and Jun Yu. Our research assistant Yuchi Liu helped us with various tasks for the empirical analysis. Leng’s research is partially supported by the National Natural Science Foundation of China Grant 72001183. Sun acknowledges the support from the National Natural Science Foundation of China under Grant Number 72203032.
@@ A Proofs
In this appendix, we prove the asymptotic properties of $L(\theta)$, the corrected likelihood, $\widehat{\theta}_{L}$ the corrected estimator, and $\{\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}\}$, the three test statistics we consider in this paper.
To continue, we first introduce three Lemmas. These lemmas are useful in establishing Theorems 1 (asymptotic bias expansion) and 2 (corrected likelihood). The proof of Corollaries 1 (estimators) and 2 (tests) then follow.
Define the partition
$H(\theta)=:\begin{pmatrix}H_{\alpha\alpha^{\prime}}\left(\theta\right),&H_{\alpha\gamma^{\prime}}\left(\theta\right)\\ H_{\gamma\alpha^{\prime}}\left(\theta\right),&H_{\gamma\gamma^{\prime}}\left(\theta\right)\end{pmatrix},$
where $H_{\alpha\alpha^{\prime}}\left(\theta\right)$ is $\left(N-1\right)K\times\left(N-1\right)K$, $H_{\gamma\gamma^{\prime}}\left(\theta\right)$ is $\left(T-1\right)K\times\left(T-1\right)K$, $H_{\alpha\gamma^{\prime}}\left(\theta\right)$ is $\left(N-1\right)K\times\left(T-1\right)K$, and $H_{\gamma\alpha^{\prime}}\left(\theta\right)=\left[H_{\alpha\gamma^{\prime}}\left(\theta\right)\right]^{\prime}$. The next lemma states that the quadratic form $\mathbb{E}\left\{s^{\prime}(\theta)[{H}\left(\theta\right)]^{-1}s(\theta)\right\}$ can be decomposed into two dominating terms of order $O(1/\sqrt{NT})$. They establish the two additive bias terms $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$, arising from estimating individual effects and time effects, respectively.
Lemma A.1
Under Assumption 1, we have
$\mathbb{E}\left\{s^{\prime}(\theta)[{H}\left(\theta\right)]^{-1}s(\theta)\right\}=\mathbb{E}\left\{\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right\}+\mathbb{E\{}\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\}+o(1/\sqrt{NT}),$
where $o(\cdot)$ is uniform in $\theta$, and
$\sup_{\theta\in\Theta}\mathbb{E}\left\{\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right\}=O\left(1/T\right),\qquad\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right]=O(1/N).$
Proof. Seeing $\mathbb{E}s(\theta)=0$, the claim without $\mathbb{E}$ is
$\quad\quad s^{\prime}(\theta)[-{H}(\theta)]^{-1}s\left(\theta\right)$
@@
$=-\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)+2\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right](1+o_{\mathbb{P}}(1)).$ (A.1)
Since $K$ is fixed, we assume without loss of generality that $K=1$. In what follows, we show
$\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right]=O\left(1/(NT)\right),$ (A.2)
@@
$\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right]=O(1/T),\qquad\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right]=O(1/N).$ (A.3)